OpenStudy (sloppycanada):

The sum of three numbers is 11. If you triple the second number and add the third number, the result will be equal to the first number. The first number is two times the second number. What are the numbers? 1st number = 9, 2nd number = 6, 3rd number = -4 1st number = 10, 2nd number = 5, 3rd number = 4 1st number = 10, 2nd number = 5, 3rd number = 14 1st number = 11, 2nd number = 5.5, 3rd number = -5.5

rishavraj (rishavraj):

let the nos be x , y , z x + y +z = 11 3y + z = x x = 2y

OpenStudy (sloppycanada):

So the second one?

rishavraj (rishavraj):

nah nah.....see as x = 2y so 3y + z = x 3y + z = 2x y + z = 0 x + y + z = 11 so as y + z = 0 z = 11

OpenStudy (sloppycanada):

Ohhh okay so it's the last one...

rishavraj (rishavraj):

hmmm yeah....!!